find three consecutive odd integers whose sum is 24 more than the greateset of the three integers
Answer by oberobic(2304)

We can define the consecutive odd integers as:
x
x+2
x+4
.
3x+6 = x+4 + 24
.
3x +6 = x +28
.
2x = 22
.
x = 11
x+2 = 13
x+4 = 15
.
Check the answer.
sum = 39
39 = 15+24
Correct
.
Answer: The three numbers are 11, 13, and 15.
.
Done.